A) Solve the following linear program graphically. How many extreme points exist for this problem?

Min 150X + 210Y
s.t. 3.8X + 1.2Y ≥ 22.8
                 Y ≥ 6
                 Y ≤ 15
45X +  30Y = 630
            X, Y ≥ 0

B)

Comfort Plus Inc. (CPI) manufactures a standard dining chair used in restaurants. The demand forecasts for quarter 1 (January–March) and quarter 2 (April–June) are 3700 chairs and 4200 chairs, respectively. CPI has a policy of satisfying all demand in the quarter in which it occurs.

The chair contains an upholstered seat that can be produced by CPI or purchased from DAP, a subcontractor.  DAP currently charges $12.50 per seat, but has announced a new price of $13.75, effective April 1.  CPI can produce the seat at a cost of $10.25.   CPI can produce up to 3800 seats per quarter.

Seats that are produced or purchased in quarter 1 and used to satisfy demand in quarter 2 cost CPI $1.50 each to hold in inventory, but maximum inventory cannot exceed 300 seats.

Formulate this problem as a linear programming problem.
Answer #1

A min Z=150x+ 210X st 3-6X + 1-27 > 22-9 YS15 45 X+307 =530 X/Y > To draw constraint 3.8*+1:27>22:8 Treat it as 3.8x+12y= 228Extieme points 0 4915) O (10) obj function value 15004+ 210 (15)= 3750 150610 † 210C6)= 2760 and mis value is occurrs at (10,